# Relations Between the Inequality Indices Gini, Pietra and Kolkata: Theory and Data Analysis

### v1

@@ Abstract
We study here relations between three inequality indices, namely the Gini ($g$), Pietra ($p$) and Kolkata ($k$) introduced in 1912, 1915 and 2014 respectively and all are derived from the Lorenz function $L(x)$ introduced in 1905. The Kolkata index (which corresponds to a fixed point of the complementary Lorenz function $L_{c}(x)\equiv 1-L(x)$) gives the fraction of wealth $k$ possessed by the richest 1k1-k fraction of people ($k$ = 0.8 corresponds to Pareto’s 80-20 law from 1896). We show rigorously that the Pietra or Robin Hood index $p$ should equal to the excess wealth fraction 2k12k-1 possessed by the richest 1k1-k fraction of people. Our numerical data analysis for US IRS Income data (1983-2022), Bollywood (India) movie income data (1999-2024) and the citation inequalities across the publications by forty Nobel Laureates (2020-2025) in Economics, Physics, Chemistry and Medicine clearly shows that p/(2k1)p/(2k-1) is always greater than unity but the deviation is never more than five percent. Assuming some simple analytic form for the Lorenz function, we also derived the relations $k=(1/2)+(3/8)g$ for small $g$ values and $p/g=3/4$. However, these relations generally deviate significantly for larger $g$ or $k$ values when compared with observations.
@@ I Introduction
Figure 2: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for US economy income data (IRS data irs ; Yakovenko22 for the period 1982-2022). The values of $k$ and $g$ are growing with time, because of increasing rate of withdrawal of public welfare programs. The upper left inset, showing the values of p/(2k1)p/(2k-1) against the years, indicate a value higher than unity as predicted by a rigid theoretical argument (relation (1)). The lower right inset shows the values of $p/g$ against years and show values considerably different from 3/4, as obtained from the theoretical relation (3) (obtained an additional assumption (2) of the minimal form of the Lorenz function). Details of the values of the inequality indices and and of their relations are given in Table I of the Appendix.
@@ II Theoretical analysis of Inequality indices
We will first explore analytically if there are any relations among these three indices, namely Gini ($g$), Pietra ($p$) and Kolkata ($k$). To begin with, we note that since by construction the Kolkata index ($k$) is given by the fraction of the total wealth possessed by the 1k1-k fraction of richest people ($k>1/2$; $k=1/2$ corresponds to equality), they possess precisely an extra amount of 2k12k-1 fraction of total wealth. The Pietra index Allen22 (or Robin Hood index) $p$ then should equal to this excess amount, suggesting
p=2k1.p=2k-1. (1)
In order to find theoretically the other plausible relations among those indices, let us proceed with a (Landau-like; see e.g., Manna22 ) minimal polynomial expansion of the Lorenz function:
@@
Also noting that the Kolkata index $k$ is given by the fixed point equation $L_{c}(k)=k$ or the complementary Lorenz function $L_{c}(x)\equiv 1-L(x)$, giving (see also Manna22 ; Banerjee23 )
@@
Figure 4: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for movie income inequalities (Bolloywood India data 13bbollywood for the period 1999-2024). The values of $k$ in the tax data (which can be argued to represent the prevailing inequality status better) seems to have grown a little beyond the Pareto value ($k$ = 0.80) with increasing withdrawal of public welfare programs. The upper left inset, showing the values of p/(2k1)p/(2k-1) against the years, indicate values higher than unity as predicted by a rigid theoretical argument (relation (1)). The lower right inset shows the values of $p/g$ against years and show values considerably different from 3/4, as obtained from theoretical relation (3) (obtained an additional assumption (2) of the minimal form of the Lorenz function). Details of the values of the inequality indices and and of their relations are given in Table III of the Appendix.
@@ IV Summary and Conclusion
As we discussed, all the three inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) are derived from the Lorenz function $L(x)$ (see Fig. 1). $2g$ is given by the area between the equality line and the Lorenz curve. $p$ is given by the maximum value of xL(x)x-L(x) and is interpreted as a measure of the fraction of money the ‘rich’ people should transfer to the ‘poor’ people (neither rich or poor are specifically defined here), so that a redistribution would bring equality (Robin Hood index Allen22 ). $k$ is given by the fixed point of the Complementary Lorenz Function $L_{c}(x)\equiv 1-L(x)$: $L_{c}(k)=k$. Geometrically (see Fig. 1) it gives the fraction $k$ of wealth (citations) possessed by clearly defined 1k1-k fraction of rich people (papers). We therefore argued the excess wealth of the rich is precisely equal to 2k12k-1, which if they transfer to the poor $k$ fraction of people, then (on redistribution) equality would be achieved. As such p=2k1p=2k-1 (relation (1)).
@@
As we see from numerical data analysis (see Appendices I-IV) and upper-left insets of Figs. 2 - 5, the $p$ is a little higher (at most $5\%$ off) from that suggested by relation (1). Next we assumed a minimal analytic and polynomial form (2) for the Lorenz function $L(x)$ and calculated $g$ and $p$ from there, giving $p=(3/4)g$ (relation (3)). As we see from the tables I-IV and Figs. 2-5, there are large deviations from relation (2). Finally, solving for $k$ from (2) and (3), in the small $g$ limit, we got $k=(1/2)+(3/8)g$ (relation (3)). Numerical data analysis for the validity of this relation are indicated in each of the main parts and bottom right insets of Figs. 2-5.
@@
We studied here relations between three inequality indices, namely the Gini ($g$) Gini1912 , Pietra ($p$) Pietra15 and Kolkata ($k$) Ghosh14 ; Banerjee20 ; Banerjee23 introduced in 1912, 1915 and 2014 respectively and all are derived from the Lorenz function Lorenz1905 $L(x)$ introduced in 1905. We showed rigorously that the Pietra or Robin Hood index $p$ should equal to the excess wealth 2k12k-1 possessed by the richest 1k1-k people, giving relation (1). Our numerical data analysis results for US IRS Income data (1983-2022; see Tables I, II of the Appendix, and Figs. 2,3), Bollywood (India) movie income data (1999-2024; see Table III of the Appendix and Fig. 4) and the citation inequalities across the publications by forty Nobel Laureates (2020-2025; see Table IV of the Appendix and Fig. 5) in Economics, Physics, Chemistry and Medicine clearly shows that p/(2k1)p/(2k-1) is always slightly greater than unity and the deviation is never more than five percent. Assuming some simple analytic form (2) for the Lorenz function (cf. Joseph22 ), we also derived here the relations (3) $p/g=3/4$ and (4) $k=(1/2)+(3/8)g$ for small values of $g$. However, these relations generally deviate significantly for larger $g$ and $k$ values (see Tables I-IV and Figs. 2-5) when compared with observations.
Acknowledgement
We are grateful to Suchismita Banerjee and Shohini Sen for their valuable comments and suggestions.
@@ Acknowledgement
Table 1: Gini-index ($g$), Pietra-index ($p$) and Kolkata-index ($k$) values for the income inequalities from movie for U.S. IRS Income Data (1983–2022) datasets collected from ref. irs ; Yakovenko22 . For each year, the total income collections of individuals are analyzed.
Year Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g 1983 0.47418 0.34280 0.66894 1.0146 0.7229 0.7126 1984 0.47711 0.34289 0.67076 1.0040 0.7187 0.7158 1985 0.48136 0.34672 0.67253 1.0048 0.7203 0.7168 1986 0.49412 0.35602 0.67640 1.0091 0.7205 0.7140 1987 0.51365 0.36991 0.68403 1.0050 0.7202 0.7166 1988 0.53394 0.38312 0.69106 1.0026 0.7175 0.7157 1989 0.53094 0.38250 0.68970 1.0082 0.7204 0.7146 1990 0.52898 0.38110 0.68854 1.0107 0.7204 0.7128 1991 0.52472 0.37906 0.68734 1.0117 0.7224 0.7141 1992 0.53185 0.38256 0.68960 1.0089 0.7193 0.7130 1993 0.53100 0.38191 0.68960 1.0071 0.7192 0.7141 1994 0.53198 0.38212 0.69071 1.0018 0.7183 0.7170 1995 0.54026 0.38946 0.69434 1.0020 0.7209 0.7194 1996 0.55188 0.39894 0.69943 1.0002 0.7229 0.7227 1997 0.55998 0.40463 0.70153 1.0039 0.7226 0.7198 1998 0.56631 0.40807 0.70376 1.0013 0.7206 0.7196 1999 0.57604 0.41659 0.70780 1.0024 0.7232 0.7215 2000 0.58418 0.42321 0.71140 1.0010 0.7245 0.7237 2001 0.56116 0.40616 0.70285 1.0011 0.7238 0.7230 2002 0.55009 0.39857 0.69902 1.0013 0.7246 0.7236 2003 0.55508 0.40325 0.70161 1.0001 0.7265 0.7264 2004 0.57073 0.41385 0.70533 1.0078 0.7251 0.7195 2005 0.58713 0.42413 0.71052 1.0073 0.7224 0.7171 2006 0.59445 0.42721 0.71290 1.0033 0.7187 0.7163 2007 0.60078 0.43191 0.71543 1.0024 0.7189 0.7172 2008 0.58395 0.42027 0.70950 1.0030 0.7197 0.7175 2009 0.56321 0.40960 0.70280 1.0099 0.7273 0.7202 2010 0.57349 0.41558 0.70664 1.0056 0.7247 0.7206 2011 0.57799 0.41882 0.70862 1.0038 0.7246 0.7219 2012 0.59759 0.43488 0.71601 1.0066 0.7277 0.7229 2013 0.58345 0.42418 0.71113 1.0045 0.7270 0.7237 2014 0.59374 0.43475 0.71569 1.0078 0.7322 0.7265 2015 0.59442 0.43629 0.71649 1.0076 0.7340 0.7284 2016 0.58799 0.43158 0.71424 1.0072 0.7340 0.7287 2017 0.59553 0.43759 0.71800 1.0036 0.7348 0.7321 2018 0.59473 0.43703 0.71829 1.0010 0.7348 0.7341 2019 0.59072 0.43407 0.71669 1.0016 0.7348 0.7336 2020 0.60100 0.44124 0.72022 1.0018 0.7342 0.7328 2021 0.62792 0.46403 0.73005 1.0085 0.7390 0.7327 2022 0.60345 0.44444 0.72064 1.0072 0.7365 0.7313
@@
Year Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g 1983 0.66335 0.48721 0.74086 1.0114 0.7345 0.7262 1984 0.66753 0.48835 0.74363 1.0022 0.7316 0.7299 1985 0.67246 0.49438 0.74581 1.0056 0.7352 0.7311 1986 0.69308 0.51074 0.75262 1.0109 0.7369 0.7290 1987 0.70073 0.51685 0.75674 1.0066 0.7376 0.7328 1988 0.71296 0.52519 0.76215 1.0017 0.7366 0.7354 1989 0.70412 0.51864 0.75899 1.0013 0.7366 0.7356 1990 0.70247 0.51917 0.75930 1.0011 0.7391 0.7383 1991 0.70487 0.52181 0.76089 1.0001 0.7403 0.7402 1992 0.71938 0.53487 0.76601 1.0054 0.7435 0.7396 1993 0.72833 0.54241 0.76992 1.0048 0.7447 0.7412 1994 0.72995 0.54421 0.77123 1.0032 0.7455 0.7431 1995 0.73822 0.55293 0.77611 1.0013 0.7490 0.7480 1996 0.74976 0.56455 0.78198 1.0010 0.7530 0.7522 1997 0.75437 0.56893 0.78243 1.0072 0.7542 0.7488 1998 0.76372 0.57613 0.78651 1.0054 0.7544 0.7503 1999 0.77301 0.58454 0.79122 1.0036 0.7562 0.7535 2000 0.77856 0.59549 0.79470 1.0103 0.7649 0.7570 2001 0.76491 0.57785 0.78725 1.0058 0.7554 0.7511 2002 0.77313 0.58737 0.79221 1.0050 0.7597 0.7559 2003 0.77373 0.58709 0.79204 1.0052 0.7588 0.7549 2004 0.78684 0.60559 0.80048 1.0077 0.7696 0.7638 2005 0.79995 0.62076 0.80998 1.0013 0.7760 0.7750 2006 0.80278 0.62305 0.81109 1.0014 0.7761 0.7750 2007 0.80582 0.62481 0.81229 1.0004 0.7754 0.7751 2008 0.80036 0.62046 0.80949 1.0024 0.7752 0.7734 2009 0.80642 0.63018 0.81371 1.0044 0.7815 0.7780 2010 0.81101 0.63382 0.81576 1.0036 0.7815 0.7787 2011 0.79568 0.61789 0.80758 1.0044 0.7766 0.7731 2012 0.80744 0.62826 0.81359 1.0017 0.7781 0.7768 2013 0.80558 0.62664 0.81260 1.0023 0.7779 0.7761 2014 0.81165 0.63613 0.81674 1.0042 0.7837 0.7805 2015 0.80929 0.63505 0.81648 1.0033 0.7847 0.7821 2016 0.80136 0.62672 0.81236 1.0032 0.7821 0.7796 2017 0.80401 0.63025 0.81495 1.0006 0.7839 0.7834 2018 0.81284 0.64080 0.81722 1.0100 0.7883 0.7805 2019 0.80787 0.63630 0.81471 1.0109 0.7876 0.7791 2020 0.82292 0.65008 0.82091 1.0129 0.7900 0.7799 2021 0.82025 0.64352 0.81709 1.0147 0.7845 0.7732 2022 0.81013 0.62775 0.81029 1.0116 0.7749 0.7660
@@
Year Tot. No. of movies Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g 1999 52 0.41093 0.29326 0.64469 1.0134 0.7136 0.7042 2000 51 0.41637 0.29554 0.64453 1.0224 0.7098 0.6942 2001 51 0.55679 0.41486 0.70681 1.0030 0.7451 0.7429 2002 51 0.36364 0.26156 0.63000 1.0060 0.7193 0.7150 2003 52 0.45222 0.33932 0.66963 1.0002 0.7503 0.7502 2004 51 0.44658 0.34524 0.67206 1.0033 0.7731 0.7706 2005 52 0.38679 0.29895 0.64882 1.0044 0.7729 0.7695 2006 50 0.52425 0.41003 0.70370 1.0065 0.7821 0.7771 2007 51 0.49243 0.36600 0.68287 1.0007 0.7433 0.7427 2008 51 0.49477 0.36377 0.68170 1.0010 0.7352 0.7345 2009 64 0.61159 0.46545 0.72633 1.0283 0.7610 0.7401 2010 140 0.77307 0.63066 0.80724 1.0263 0.8158 0.7949 2011 123 0.78239 0.65696 0.81366 1.0472 0.8397 0.8018 2012 133 0.78140 0.64864 0.81088 1.0432 0.8301 0.7957 2013 137 0.76403 0.60528 0.79360 1.0308 0.7922 0.7686 2014 146 0.80382 0.65775 0.81760 1.0355 0.8183 0.7902 2015 165 0.79784 0.65836 0.81791 1.0355 0.8252 0.7969 2016 216 0.82970 0.68791 0.83025 1.0415 0.8291 0.7961 2017 252 0.85300 0.70442 0.84432 1.0229 0.8258 0.8073 2018 219 0.84407 0.71848 0.84777 1.0330 0.8512 0.8240 2019 244 0.84798 0.72439 0.85467 1.0212 0.8543 0.8365 2020 76 0.86551 0.73448 0.85860 1.0241 0.8486 0.8286 2021 50 0.79775 0.63359 0.81054 1.0201 0.7942 0.7785 2022 165 0.84379 0.69608 0.84151 1.0191 0.8249 0.8095 2023 227 0.89157 0.77006 0.87826 1.0179 0.8637 0.8485 2024 230 0.89728 0.76409 0.87516 1.0184 0.8516 0.8362

### v2

@@ Abstract
We study here relations between three inequality indices, namely the Gini ($g$), Pietra ($p$) and Kolkata ($k$) introduced in 1912, 1915 and 2014 respectively and all are derived from the Lorenz function $L(x)$ introduced in 1905. The Kolkata index (which corresponds to a fixed point of the complementary Lorenz function $L_{c}(x)\equiv 1-L(x)$) gives the fraction of wealth $k$ possessed by the richest 1k1-k fraction of people ($k$ = 0.8 corresponds to Pareto’s 80-20 law from 1896). We show rigorously that while the Pietra index value $p$ should be greater than or equal to 2k12k-1, the Robin Hood index should strictly be equal to the excess wealth fraction 2k12k-1 possessed by the richest 1k1-k fraction of people. Our numerical data analysis for US IRS Income data (1983-2022), Bollywood (India) movie income data (1999-2024) and the citation inequalities across the publications by forty Nobel Laureates (2020-2025) in Economics, Physics, Chemistry and Medicine clearly shows that p/(2k1)p/(2k-1) is always greater than unity but the deviation is never more than five percent. Assuming some simple analytic form for the Lorenz function, we also derived the relations $k=(1/2)+(3/8)g$ for small $g$ values and $p/g=3/4$. However, these relations generally deviate significantly for larger $g$ or $k$ values when compared with observations.
@@ I Introduction
Figure 2: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for US economy income data (IRS data irs ; Yakovenko22 for the period 1982-2022). The values of $k$ and $g$ are growing with time, because of increasing rate of withdrawal of public welfare programs. The upper left inset, showing the values of p/(2k1)p/(2k-1) against the years, indicate a value higher than unity as predicted by a rigid theoretical argument (relation (3). The lower right inset shows the values of $p/g$ against years and show values considerably different from 3/4, as obtained from the theoretical relation (5) (obtained an additional assumption (4) of the minimal form of the Lorenz function). Details of the values of the inequality indices and and of their relations are given in Table I of the Appendix.
@@ II Theoretical analysis of Inequality indices
We will first explore analytically if there are any relations among these three indices, namely Gini ($g$), Pietra ($p$) and Kolkata ($k$). To begin with, we note that since by construction the Kolkata index ($k$) is given by the fraction of the total wealth possessed by the 1k1-k fraction of richest people ($k>1/2$; $k=1/2$ corresponds to equality), they possess precisely an extra amount of 2k12k-1 fraction of total wealth. The Pietra index $p$ , if strictly interpreted as the Robin Hood index Allen22 , should equal to this excess amount equal to 2k12k-1.
On the other hand, the Pietra index $p$ is defined as the maximum vertical distance between the Lorenz curve $L(x)$ and the line of perfect equality:
$p=\max_{x}\left(x-L(x)\right).$ (1)
Evaluating the vertical distance between the equality line and the Lorenz curve at $x=k$, we obtain
$\left.x-L(x)\right|_{x=k}=k-(1-k)=2k-1.$ (2)
Since the Pietra index is the maximum value of this vertical distance over all $x$, it follows that
$p\geq 2k-1.$ (3)
In order to find theoretically the other plausible relations among those indices, let us proceed with a (Landau-like; see e.g., Joseph22 ) minimal polynomial expansion of the Lorenz function:
@@
Also noting that the Kolkata index $k$ is given by the fixed point equation $L_{c}(k)=k$ or the complementary Lorenz function $L_{c}(x)\equiv 1-L(x)$, giving (see also Joseph22 ; Banerjee23B )
@@
Figure 4: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for movie income inequalities (Bolloywood India data 13bbollywood for the period 1999-2024). The values of $k$ in the tax data (which can be argued to represent the prevailing inequality status better) seems to have grown a little beyond the Pareto value ($k$ = 0.80) with increasing withdrawal of public welfare programs. The upper left inset, showing the values of p/(2k1)p/(2k-1) against the years, indicate values higher than unity as predicted by a rigid theoretical argument (relation (3). The lower right inset shows the values of $p/g$ against years and show values considerably different from 3/4, as obtained from theoretical relation (5) (obtained an additional assumption (4) of the minimal form of the Lorenz function). Details of the values of the inequality indices and and of their relations are given in Table III of the Appendix.
@@ IV Summary and Conclusion
As we discussed, all the three inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) are derived from the Lorenz function $L(x)$ (see Fig. 1). $2g$ is given by the area between the equality line and the Lorenz curve. $p$ is given by the maximum value of xL(x)x-L(x) and is interpreted as a measure of the fraction of money the ‘rich’ people should transfer to the ‘poor’ people (neither rich or poor are specifically defined here), so that a redistribution would bring equality (Robin Hood index Allen22 ). $k$ is given by the fixed point of the Complementary Lorenz Function $L_{c}(x)\equiv 1-L(x)$: $L_{c}(k)=k$. Geometrically (see Fig. 1) it gives the fraction $k$ of wealth (citations) possessed by clearly defined 1k1-k fraction of rich people (papers). We therefore argued the excess wealth of the rich is precisely equal to 2k12k-1, which if they transfer to the poor $k$ fraction of people, then (on redistribution) equality would be achieved. As such, the Robin Hood index should be equal to 2k12k-1, while the value of Pietra Index $p$, given by the maximum vertical distance of the Lorenz curve from the equality line is given by the relation (3), with 2k12k-1 as its minimum value.
@@
As we see from numerical data analysis (see Appendices I-IV) and upper-left insets of Figs. 2 - 5, the $p$ is a little higher (at most 5% off) from the value (2k12k-1 of the Robin Hood index), as given by the relation (3) . Next we assumed a minimal analytic and polynomial form 4 for the Lorenz function $L(x)$ and calculated $g$ and $p$ from there, giving $p=(3/4)g$ (relation (5)). As we see from the tables I-IV and Figs. 2-5, there are large deviations from relation (5). Finally, solving for $k$ from (4) and (5), in the small $g$ limit, we got $k=(1/2)+(3/8)g$ (relation (6)). Numerical data analysis for the validity of this relation are indicated in each of the main parts and bottom right insets of Figs. 2-5.
@@
We studied here relations between three inequality indices, namely the Gini ($g$) Gini1912 , Pietra ($p$) Pietra15 and Kolkata ($k$) Ghosh14 ; Banerjee20 ; Banerjee23 introduced in 1912, 1915 and 2014 respectively and all are derived from the Lorenz function Lorenz1905 $L(x)$ introduced in 1905. We showed rigorously that while the Pietra index $p$ should have the minimum value 2k12k-1 (relation (3), the Robin Hood index value will be given the faction 2k12k-1 of excess wealth possessed by the richest 1k1-k fraction of people. Our numerical data analysis results for US IRS Income data (1983-2022; see Tables I, II of the Appendix, and Figs. 2,3), Bollywood (India) movie income data (1999-2024; see Table III of the Appendix and Fig. 4) and the citation inequalities across the publications by forty Nobel Laureates (2020-2025; see Table IV of the Appendix and Fig. 5) in Economics, Physics, Chemistry and Medicine clearly shows that p/(2k1)p/(2k-1) is always slightly greater than unity and the deviation is never more than five percent. Assuming some simple analytic form (4) for the Lorenz function (cf. Joseph22 ), we also derived here the relations (5) $p/g=3/4$ and (6) $k=(1/2)+(3/8)g$ for small values of $g$. However, these relations generally deviate significantly for larger $g$ and $k$ values (see Tables I-IV and Figs. 2-5) when compared with observations.
@@
Table 1: Gini-index ($g$), Pietra-index ($p$) and Kolkata-index ($k$) values for the income inequalities from U.S. IRS Income Data (1983–2022) collected from ref. irs ; Yakovenko22 . For each year, the total income collections of individuals are analyzed.
Year
Gini index ($g$)
Pietra index ($p$)
Kolkata index ($k$)
p/(2k1)p/(2k-1)
$p/g$
(2k1)/g(2k-1)/g
1983 0.47418 0.34280 0.66894 1.0146 0.7229 0.7126
1984 0.47711 0.34289 0.67076 1.0040 0.7187 0.7158
1985 0.48136 0.34672 0.67253 1.0048 0.7203 0.7168
1986 0.49412 0.35602 0.67640 1.0091 0.7205 0.7140
1987 0.51365 0.36991 0.68403 1.0050 0.7202 0.7166
1988 0.53394 0.38312 0.69106 1.0026 0.7175 0.7157
1989 0.53094 0.38250 0.68970 1.0082 0.7204 0.7146
1990 0.52898 0.38110 0.68854 1.0107 0.7204 0.7128
1991 0.52472 0.37906 0.68734 1.0117 0.7224 0.7141
1992 0.53185 0.38256 0.68960 1.0089 0.7193 0.7130
1993 0.53100 0.38191 0.68960 1.0071 0.7192 0.7141
1994 0.53198 0.38212 0.69071 1.0018 0.7183 0.7170
1995 0.54026 0.38946 0.69434 1.0020 0.7209 0.7194
1996 0.55188 0.39894 0.69943 1.0002 0.7229 0.7227
1997 0.55998 0.40463 0.70153 1.0039 0.7226 0.7198
1998 0.56631 0.40807 0.70376 1.0013 0.7206 0.7196
1999 0.57604 0.41659 0.70780 1.0024 0.7232 0.7215
2000 0.58418 0.42321 0.71140 1.0010 0.7245 0.7237
2001 0.56116 0.40616 0.70285 1.0011 0.7238 0.7230
2002 0.55009 0.39857 0.69902 1.0013 0.7246 0.7236
2003 0.55508 0.40325 0.70161 1.0001 0.7265 0.7264
2004 0.57073 0.41385 0.70533 1.0078 0.7251 0.7195
2005 0.58713 0.42413 0.71052 1.0073 0.7224 0.7171
2006 0.59445 0.42721 0.71290 1.0033 0.7187 0.7163
2007 0.60078 0.43191 0.71543 1.0024 0.7189 0.7172
2008 0.58395 0.42027 0.70950 1.0030 0.7197 0.7175
2009 0.56321 0.40960 0.70280 1.0099 0.7273 0.7202
2010 0.57349 0.41558 0.70664 1.0056 0.7247 0.7206
2011 0.57799 0.41882 0.70862 1.0038 0.7246 0.7219
2012 0.59759 0.43488 0.71601 1.0066 0.7277 0.7229
2013 0.58345 0.42418 0.71113 1.0045 0.7270 0.7237
2014 0.59374 0.43475 0.71569 1.0078 0.7322 0.7265
2015 0.59442 0.43629 0.71649 1.0076 0.7340 0.7284
2016 0.58799 0.43158 0.71424 1.0072 0.7340 0.7287
2017 0.59553 0.43759 0.71800 1.0036 0.7348 0.7321
2018 0.59473 0.43703 0.71829 1.0010 0.7348 0.7341
2019 0.59072 0.43407 0.71669 1.0016 0.7348 0.7336
2020 0.60100 0.44124 0.72022 1.0018 0.7342 0.7328
2021 0.62792 0.46403 0.73005 1.0085 0.7390 0.7327
2022 0.60345 0.44444 0.72064 1.0072 0.7365 0.7313
@@
Year
Gini index ($g$)
Pietra index ($p$)
Kolkata index ($k$)
p/(2k1)p/(2k-1)
$p/g$
(2k1)/g(2k-1)/g
1983 0.66335 0.48721 0.74086 1.0114 0.7345 0.7262
1984 0.66753 0.48835 0.74363 1.0022 0.7316 0.7299
1985 0.67246 0.49438 0.74581 1.0056 0.7352 0.7311
1986 0.69308 0.51074 0.75262 1.0109 0.7369 0.7290
1987 0.70073 0.51685 0.75674 1.0066 0.7376 0.7328
1988 0.71296 0.52519 0.76215 1.0017 0.7366 0.7354
1989 0.70412 0.51864 0.75899 1.0013 0.7366 0.7356
1990 0.70247 0.51917 0.75930 1.0011 0.7391 0.7383
1991 0.70487 0.52181 0.76089 1.0001 0.7403 0.7402
1992 0.71938 0.53487 0.76601 1.0054 0.7435 0.7396
1993 0.72833 0.54241 0.76992 1.0048 0.7447 0.7412
1994 0.72995 0.54421 0.77123 1.0032 0.7455 0.7431
1995 0.73822 0.55293 0.77611 1.0013 0.7490 0.7480
1996 0.74976 0.56455 0.78198 1.0010 0.7530 0.7522
1997 0.75437 0.56893 0.78243 1.0072 0.7542 0.7488
1998 0.76372 0.57613 0.78651 1.0054 0.7544 0.7503
1999 0.77301 0.58454 0.79122 1.0036 0.7562 0.7535
2000 0.77856 0.59549 0.79470 1.0103 0.7649 0.7570
2001 0.76491 0.57785 0.78725 1.0058 0.7554 0.7511
2002 0.77313 0.58737 0.79221 1.0050 0.7597 0.7559
2003 0.77373 0.58709 0.79204 1.0052 0.7588 0.7549
2004 0.78684 0.60559 0.80048 1.0077 0.7696 0.7638
2005 0.79995 0.62076 0.80998 1.0013 0.7760 0.7750
2006 0.80278 0.62305 0.81109 1.0014 0.7761 0.7750
2007 0.80582 0.62481 0.81229 1.0004 0.7754 0.7751
2008 0.80036 0.62046 0.80949 1.0024 0.7752 0.7734
2009 0.80642 0.63018 0.81371 1.0044 0.7815 0.7780
2010 0.81101 0.63382 0.81576 1.0036 0.7815 0.7787
2011 0.79568 0.61789 0.80758 1.0044 0.7766 0.7731
2012 0.80744 0.62826 0.81359 1.0017 0.7781 0.7768
2013 0.80558 0.62664 0.81260 1.0023 0.7779 0.7761
2014 0.81165 0.63613 0.81674 1.0042 0.7837 0.7805
2015 0.80929 0.63505 0.81648 1.0033 0.7847 0.7821
2016 0.80136 0.62672 0.81236 1.0032 0.7821 0.7796
2017 0.80401 0.63025 0.81495 1.0006 0.7839 0.7834
2018 0.81284 0.64080 0.81722 1.0100 0.7883 0.7805
2019 0.80787 0.63630 0.81471 1.0109 0.7876 0.7791
2020 0.82292 0.65008 0.82091 1.0129 0.7900 0.7799
2021 0.82025 0.64352 0.81709 1.0147 0.7845 0.7732
2022 0.81013 0.62775 0.81029 1.0116 0.7749 0.7660
@@
Year
Tot. No. of movies
Gini index ($g$)
Pietra index ($p$)
Kolkata index ($k$)
p/(2k1)p/(2k-1)
$p/g$
(2k1)/g(2k-1)/g
1999 52 0.41093 0.29326 0.64469 1.0134 0.7136 0.7042
2000 51 0.41637 0.29554 0.64453 1.0224 0.7098 0.6942
2001 51 0.55679 0.41486 0.70681 1.0030 0.7451 0.7429
2002 51 0.36364 0.26156 0.63000 1.0060 0.7193 0.7150
2003 52 0.45222 0.33932 0.66963 1.0002 0.7503 0.7502
2004 51 0.44658 0.34524 0.67206 1.0033 0.7731 0.7706
2005 52 0.38679 0.29895 0.64882 1.0044 0.7729 0.7695
2006 50 0.52425 0.41003 0.70370 1.0065 0.7821 0.7771
2007 51 0.49243 0.36600 0.68287 1.0007 0.7433 0.7427
2008 51 0.49477 0.36377 0.68170 1.0010 0.7352 0.7345
2009 64 0.61159 0.46545 0.72633 1.0283 0.7610 0.7401
2010 140 0.77307 0.63066 0.80724 1.0263 0.8158 0.7949
2011 123 0.78239 0.65696 0.81366 1.0472 0.8397 0.8018
2012 133 0.78140 0.64864 0.81088 1.0432 0.8301 0.7957
2013 137 0.76403 0.60528 0.79360 1.0308 0.7922 0.7686
2014 146 0.80382 0.65775 0.81760 1.0355 0.8183 0.7902
2015 165 0.79784 0.65836 0.81791 1.0355 0.8252 0.7969
2016 216 0.82970 0.68791 0.83025 1.0415 0.8291 0.7961
2017 252 0.85300 0.70442 0.84432 1.0229 0.8258 0.8073
2018 219 0.84407 0.71848 0.84777 1.0330 0.8512 0.8240
2019 244 0.84798 0.72439 0.85467 1.0212 0.8543 0.8365
2020 76 0.86551 0.73448 0.85860 1.0241 0.8486 0.8286
2021 50 0.79775 0.63359 0.81054 1.0201 0.7942 0.7785
2022 165 0.84379 0.69608 0.84151 1.0191 0.8249 0.8095
2023 227 0.89157 0.77006 0.87826 1.0179 0.8637 0.8485
2024 230 0.89728 0.76409 0.87516 1.0184 0.8516 0.8362